Consider a sample of nitrogen gas at volume 2728 mL at an unknown temperature. When the sample is submerged in ice water (T = 0°C), its volume changes to 1.7 L. What was the initial temperature of the nitrogen (in °C)?


1
Expert's answer
2020-09-29T06:46:00-0400

"\frac{V1}{T1}=\frac{V2}{T2}"


"\frac{2.728}{T1}=\frac{1.7}{273}"


"T1=2.728:\frac{1.7}{273}=440 K"